OpenStudy (anonymous):

Solve: 4y2 + 12y + 9 = 0

OpenStudy (anonymous):

that answer is showin up

jimthompson5910 (jim_thompson5910):

\[\Large y = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large y = \frac{-(12)\pm\sqrt{(12)^2-4(4)(9)}}{2(4)}\] \[\Large y = \frac{-12\pm\sqrt{144-(144)}}{8}\] \[\Large y = \frac{-12\pm\sqrt{0}}{8}\] \[\Large y = \frac{-12+\sqrt{0}}{8} \ \text{or} \ y = \frac{-12-\sqrt{0}}{8}\] \[\Large y = \frac{-12+0}{8} \ \text{or} \ y = \frac{-12-0}{8}\] \[\Large y = \frac{-12}{8} \ \text{or} \ y = \frac{-12}{8}\] \[\Large y = -\frac{3}{2} \ \text{or} \ y = -\frac{3}{2}\] So the only solution is \[\Large y = -\frac{3}{2}\]

OpenStudy (hoblos):

you can also get benefit of some special formulas... we know that a^2 + 2ab + b^2 =(a+b)^2 so 4y2 + 12y + 9 = 0 (2y)^2 +2(y)(3) + 3^2 =0 (2y+3)^2 =0 2y=-3 y=-3/2

hero (hero):

@hoblos solution is much more impressive